### v2

# Mingchen Li1, Jiatan Huang1, Jeremy Yeung1, Anne Blaes2 Steven Johnson3, Hongfang Liu4, Hua Xu5, Rui Zhang1

Abstract
Medical Large Language Models (LLMs) such as ClinicalCamel 70B, Llama3-OpenBioLLM 70B have demonstrated impressive performance on a wide variety of medical NLP task. However, there still lacks a large language model (LLM) specifically designed for cancer domain. Moreover, these LLMs typically have billions of parameters, making them computationally expensive for healthcare systems. Thus, in this study, we propose CancerLLM, a model with 7 billion parameters and a Mistral-style architecture, pre-trained on 2,676,642 clinical notes and 515,524 pathology reports covering 17 cancer types, followed by fine-tuning on two cancer-relevant tasks, including cancer phenotypes extraction, and cancer diagnosis generation. Our evaluation demonstrated that CancerLLM achieves state-of-the-art results compared to other existing LLMs, with an average F1 score improvement of 7.61%. Additionally, CancerLLM outperforms other models on two proposed robustness testbeds. This illustrates that CancerLLM can be effectively applied to clinical AI systems, enhancing clinical research and healthcare delivery in the field of cancer.
1 Introduction
Recently, large language models (LLMs) like GPT-4111https://openai.com/index/gpt-4/, and LLama2 (Touvron et al., 2023) have become the dominant technology in various natural language processing (NLP) tasks, these models develop impressive capabilities in different NLP tasks, including information extraction (e.g., (Li & Zhang, 2023; Li et al., 2024a; c; Zhang et al., 2024)), link prediction (e.g., (Li et al., 2024b; Liu et al., 2024; Li et al., 2022)), and question answering (e.g., (Zhuang et al., 2024; Jiang et al., 2024)). While much attention has been given to these models’ capabilities in the general domain, it’s evident that specialist models have the potential to significantly help clinical and biomedical research Singhal et al. (2023). To tailor LLMs for the clinical domain, several specialized models have been developed, such as ClinicalCamel 70B (Toma et al., 2023) and Llama3-OpenBioLLM 70B [^2] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B. Improving domain-specific language models will accelerate breakthroughs in clinical and biomedical research, leading to enhanced patient care.
Despite showcasing impressive general capabilities, current medical LLMs face significant challenges when applied to cancer. This is primarily due to a deficiency in cancer-specific knowledge within these models. This lack of specialized understanding poses a barrier to the effective utilization of LLMs in assisting doctors with cancer diagnosis and treatment planning. Expanding cancer-specific knowledge within LLMs is imperative for surmounting these obstacles and harnessing their full potential in the cancer field. By integrating comprehensive information encompassing aspects like cancer diagnosis, subjective information, objective information, nursing review of systems (nursing ros) records, laboratory test results, patient self-descriptions, medical history, treatment modalities, assessments, and patient outcomes, LLMs can furnish healthcare professionals with more precise and personalized recommendations. This augmented capability holds the promise of revolutionizing cancer care and aiding medical professionals in enhancing diagnostic accuracy and formulating treatment plans. To the best of our knowledge, CancerBERT (Zhou et al., 2022) is the only model specifically designed for the cancer domain. While CancerBERT (Zhou et al., 2022) stands out as a specialized model for cancer-related tasks, its focus is primarily on breast cancer, leaving other cancer types relatively unaddressed. This limitation underscores the need for further development of specialized models that cover a broader spectrum of cancer types and applications. By expanding the scope of such models to encompass various cancer types and different applications, healthcare professionals can access more comprehensive and tailored resources to enhance patient care in the cancer domain.
Moreover, the deployment of large parameter LLMs presents a significant challenge for hospitals or medical institutions with limited computational resources. Models with a dozen billions of parameters, such as those models with 13 billion parameters, like PMC LLaMA 13B (Wu et al., 2023) and Medalpaca 13B (Han et al., 2023), or those with 70 billion parameters, such as Llama3-OpenBioLLM-70B [^3] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B and ClinicalCamel-70B (Toma et al., 2023), pose significant computational challenges for hospitals or medical institutions with limited resources, require substantial computing power and infrastructure to train and deploy effectively. The financial and technical constraints associated with running these models may pose barriers to their widespread adoption in cancer settings. As such, there is a pressing need to develop a smaller LLM for the cancer domain, reducing computational requirements to ensure that advanced LLMs can benefit a broader range of medical professionals and patients.
Figure 1: The evolution of medical LLM performance on cancer phenotype extraction, and diagnosis generation is measured using the average F1 score, which includes Exact Match, BLEU-2, and ROUGE-L. Our CancerLLM achieves the highest performance with an F1 score of 74.54%.
Bridging these gaps, in this work, we introduce CancerLLM, a state-of-the-art LLM with 7 billion parameters designed to elevate LLM’s proficiency in the cancer domain, and aid medical professionals in cancer phenotype extraction, cancer diagnosis generation. Our model represents an important step towards democratizing advancements in the cancer domain of LLMs. More specifically, our research focuses on four steps: 1) pre-training the LLM in the style of Mistral. We pre-train the Mistral 7B using 2,676,642 cancer clinical notes and 515,524 pathology reports from the University of Minnesota (UMN) Clinical Data Repository. 2) fine-tuning the LLM on a specialized dataset. We proposed three datasets focusing on phenotype extraction, and diagnosis generation. Following this, we fine-tuned CancerLLM using instruction learning for the aforementioned tasks. 3) We evaluate the generation ability of CancerLLM against other medical LLMs using Exact Match (Rajpurkar et al., 2016), BLEU-2 (Papineni et al., 2002), and ROUGE-L (Cohan & Goharian, 2016) metrics. 4) To further evaluate the model’s ability, we propose two robustness testbeds which include counterfactual robustness and misspellings robustness in the special tasks. As shown in Figure  1, our model not only outperforms existing 7B and 13B models by a significant margin but also delivers good results with those of 70B models.
The contributions of this work can be summarized as follows:
  • We proposed CancerLLM, a large language model which focuses on the cancer domain.
  • An original study with the introduction of a benchmark of phenotype extraction and diagnosis generation, facilitating assessment against existing state-of-the-art open-source medical LLMs.
  • We have developed three datasets aimed at fine-tuning CancerLLM for its two downstream tasks.
  • A large in-depth quantitative analysis of the model’s generation ability and robustness on proposed two testbeds.
2 CANCERLLM
Figure 2: Overview of CancerLLM
Figure 2 illustrates the workflow of the CancerLLM, beginning with the injection of cancer knowledge and followed by supervised instruction tuning. Subsequently, we assess the generative capability of our framework compared to current medical LLMs using three evaluation metrics. Additionally, we introduce two testbeds to evaluate the robustness of the LLMs. Finally, the trained CancerLLM is applied to three specific tasks.
2.1 Pre-training Data for Cancer
The data used in this study were obtained from the University of Minnesota (UMN) Clinical Data Repository. It contains the health records of 31,465 cancer patients. Specifically, It includes 2,676,642 cancer clinical notes (7.27GB) and 515,524 cancer pathology reports (536.85MB). We obtained the data with the approval of the UMN Institutional Review Board (IRB).
2.2 Instruction Tuning Data for Cancer
To adapt CancerLLM for various downstream applications, we employ instruction tuning for three distinct tasks: cancer phenotypes extraction, cancer diagnosis generation, and cancer treatment plan generation. This section provides detailed information on each task description and its instruction tuning data.
2.2.1 Cancer Phenotypes extraction
Task description By providing sentences from cancer pathology reports or cancer clinical notes, the model needs to extract eight specific entities: hormone receptor type, hormone receptor status, tumor size, tumor site, cancer grade, histological type, tumor laterality, and cancer stage.
Dataset The dataset (CancerNER) used in CancerBERT (Zhou et al., 2022) focuses on named entity recognition (NER) of breast cancer-related phenotypes. This data is extracted from both clinical notes and pathology reports within electronic health records (EHRs). By inputting a sentence, CancerBERT is expected to identify and recognize each entity corresponding to the aforementioned types. LLMs have been found to underperform in NER tasks compared to supervised models tailored for sequence labeling (Wang et al., 2023). This is primarily due to the inherent differences between sequence labeling tasks, which require the model to tag each token in a sequence, and the generation tasks that LLMs are typically optimized for generating the next tokens. So, in this paper, to enhance the performance of NER, we transformed the CancerNER dataset into a question-answering format by posing eight questions based on the entity types, the answers are relevant entities. The question types are shown in Table 1. Note: by giving the input sentence (context), when the question is not relevant with the context, our model is expected to output "Not relevant". . Each instance follows the template {instruction, context, response}. The instruction is: You are a medical expert, this task involves answering the question based on the provided context or text., The context is the sentence that needs to be processed, along with a question. The response is the identified entity.
question
1 What is the tumor size in the given context?
2 What is the histological type in the given context?
3 Please identify the receptors mentioned in the provided context.
4 What is the receptor type in the given context?
5 Please identify the value of tumor laterality in the provided context.
6 What is the stage of cancer in the given context?
7 Please describe the tumor location in the given context
8 What is the grade of cancer in the given context?
Table 1: Question types
2.2.2 Diagnosis Generation
Task description By giving the information from cancer clinical notes, which includes the 1) reason for visit, 2) treatment site, 3) subjective information, 4) nursing Review of Systems (ROS), 5) objective observations, and 6) laboratory test results, the model is expected to generate the correct cancer diagnosis.
Dataset We randomly selected 10,635 cancer clinical notes that do not overlap with the pre-training dataset, and split them into 80% instances for training, 10% instances for testing, and 10% instances for validation. These datasets do not overlap. Each instance follows the template {instruction, context, response}. The instruction is: You are a medical expert. This task involves generating the diagnosis based on the provided context or text, the context includes the reason for visit, treatment site, subjective information, nursing Review of Systems (ROS), objective observations, and laboratory test results. The response is the cancer diagnosis. In general, when recording a diagnosis in the clinical notes, doctors may describe the same diagnosis using different terms or abbreviations, such as ’L lung cancer’ to refer to ’left lung cancer.’ To standardize these diagnoses, ICD codes are introduced. However, mapping these varied descriptions to a unified ICD code is a challenging task. To evaluate our model’s performance, we also created a test dataset containing 374 clinical notes, where the ground truth diagnosis is represented by the ICD code stored in the database. This testing dataset is called ICDdiagnosis.
2.3 Training Methodology
Our training methods include two phases: continued pre-training and instruction tuning. In this work, we choose Mistral 7B (Jiang et al., 2023) as the foundation model to inject cancer knowledge in the pre-training progress, as it has shown superior performance on various medical benchmarks. To improve the efficiency of the training process, we utilize Low-Rank Adaptation (LoRA) (Hu et al., 2021).
2.3.1 Continued Pre-training
This stage employs Mistral 7B, focusing on next-token prediction on clinical notes and pathology reports. Due to computing resource limitations, we undergo pre-training using LoRA. We parallelize the computation across ten 80G A100 GPUs, with each A100 utilizing around 40G of GPU memory. The total training time is 47 hours. The training batch size and evaluation batch size per device are set to 4. For LoRA, the rank is set to 8, alpha is set to 16, and dropout is set to 0.05 to optimize learning. We use the AdamW optimizer and set the learning rate to 2e-4. The training epoch is 1, and the evaluation step is 0.2.
2.3.2 Instruction Tuning
After continued pre-training, our model undergoes instruction tuning with three datasets, aligning with different clinical requirements. We also use LoRA to fine-tune the model. The rank is set to 64, and alpha is set to 16. The training step is 5. We use the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 2e-4. The model was fine-tuned using a single A100 GPU with 4 hours of training time and 7338M of GPU memory.
2.4 Baselines and Evaluation Metrics
In our work, we used 14 widely used clinical, biomedical, and general LLMs as the baseline models, including: six 7B LLMs ( PMC LLaMA 7B Wu et al. (2023), Medalpaca 7B Han et al. (2023), LLAMA-2 7B Han et al. (2023), Mistral 1*7B Jiang et al. (2023), Mixtral 8*7B Jiang et al. (2023) , and Bio-Mistral 7B Labrak et al. (2024) ), one 8B LLM ( LLama3 8B [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B), four 13B models (MedLLaMA 13B Wu et al. (2023) , PMC LLaMA 13B Wu et al. (2023), Medalpaca 13B Han et al. (2023), and LLaMA2 13B Touvron et al. (2023) ) three 70B LLMs ( LLaMA2 70B Touvron et al. (2023) , Llama3-OpenBioLLM-70B [^5] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B and ClinicalCamel-70B Toma et al. (2023) ). Same with  Li & Huang (2023), we evaluate all the models based on generative evaluation metrics, including Exact Match (Rajpurkar et al., 2016) , BLEU-2 (Papineni et al., 2002), and ROUGE-L (Cohan & Goharian, 2016).
2.5 Robustness Testbeds
Counterfactual Robustness
[…]
Misspellings Robustness
[…]
3 Results
3.1 Main Results
Exact Match BLEU-2 ROUGE-L F1
Approach Precision Recall F1 Precision Recall F1 Precision Recall F1 Average
PMC LLaMA 7B 47.04 47.04 47.04 56.41 56.41 56.41 66.98 66.98 66.98 56.81
Medalpaca 7B 41.75 41.75 41.75 50.16 50.16 50.16 62.07 62.07 62.07 51.33
LLAMA-2 7B 33.18 33.18 33.18 42.80 42.80 42.80 55.09 55.09 55.09 43.96
Mistral 1*7B 45.40 45.40 45.40 54.29 54.29 54.29 65.52 65.52 65.52 55.07
Mixtral 8*7B 51.32 51.32 51.32 59.35 59.35 59.35 69.70 69.70 69.70 60.12
Bio-Mistral 7B 62.26 62.26 62.26 68.40 68.40 68.40 76.02 76.02 76.02 68.89
LLama3 8B 51.60 51.60 51.60 60.13 60.13 60.13 69.97 69.97 69.97 60.57
MedLLaMA 13B 39.02 39.02 39.02 48.98 48.98 48.98 60.63 60.63 60.63 49.54
PMC LLaMA 13B 54.97 54.97 54.97 62.56 62.56 62.56 71.06 71.06 71.06 62.86
Medalpaca 13B 40.66 40.66 40.66 50.58 50.58 50.58 62.40 62.40 62.40 51.21
LLaMA2 13B 45.76 45.76 45.76 54.70 54.70 54.70 65.80 65.80 65.80 55.42
LLaMA2 70B 50.23 50.23 50.23 59.63 59.63 59.63 69.25 69.25 69.25 59.70
Llama3-OpenBioLLM-70B 54.42 54.42 54.42 62.36 62.36 62.36 71.65 71.65 71.65 62.81
ClinicalCamel-70B 54.60 54.60 54.60 63.34 63.34 63.34 72.73 72.73 72.73 63.55
CancerLLM(Ours) 83.50 83.50 83.50 86.60 86.60 86.60 90.34 90.34 90.34 86.81
Table 2: Results of Cancer Diagnosis Generation
3.1.1 Cancer Diagnosis Generation
To assess the scalability of our model, in this part, we evaluate the performance of CancerLLM in diagnosis generation task. Table 2 presents the experiment results of various medical LLMs. We have the following observations: (1) our CancerLLM significantly outperforms all the strong baselines and its variants across all evaluation metrics. (2) We observed that CancerLLM improve the original Mistral 1*7B, and Bio-Mistral 7B by 28.93%, and 17.92% respectively, in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (3) Our model not only achieves the best performance among 7B models but also surpasses the performance of medical LLMs with 13B and 70B parameters. (4) Without domain knowledge, larger parameter LLMs are ineffective. For instance, Llama3-OpenBioLLM-70B and ClinicalCamel-70B achieve only 62.81% and 63.55%, separately. (5) Table 3 presents the experiment results of various medical LLMs on ICDdiagnosis, We observed that CancerLLM improve the best best baseline model Bio-Mistral 7B by 4.36% in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (6) The 70B models do not achieve the best performance.
Exact Match BLEU-2 ROUGE-L F1
Approach Precision Recall F1 Precision Recall F1 Precision Recall F1 Average
PMC LLaMA 7B 24.33 24.33 24.33 35.66 35.66 35.66 44.66 44.66 44.66 34.90
Medalpaca 7B 27.27 27.27 27.27 38.09 38.09 38.09 47.86 47.86 47.86 37.73
LLAMA-2 7B 24.06 24.06 24.06 35.02 35.02 35.02 44.36 44.36 44.36 34.48
Mistral 1*7B 26.74 26.74 26.74 39.00 39.00 39.00 48.53 48.53 48.53 38.09
Mixtral 8*7B 29.68 29.68 29.68 38.18 38.18 38.18 47.86 47.86 47.86 38.57
Bio-Mistral 7B 30.21 30.21 30.21 41.16 41.16 41.16 50.69 50.69 50.69 40.68
LLama3 8B 27.01 27.01 27.01 38.16 38.16 38.16 48.05 48.05 48.05 37.73
MedLLaMA 13B 29.41 29.41 29.41 39.80 39.80 39.80 48.34 48.34 48.34 39.18
PMC LLaMA 13B 26.74 26.74 26.74 37.02 37.02 37.02 46.03 46.03 46.03 36.59
Medalpaca 13B 25.13 25.13 25.13 35.57 35.57 35.57 45.31 45.31 45.31 35.33
LLaMA2 13B 22.99 22.99 22.99 34.08 34.08 34.08 43.42 43.42 43.42 33.49
LLaMA2 70B 27.81 27.81 27.81 38.85 38.85 38.85 47.94 47.94 47.94 38.20
Llama3-OpenBioLLM-70B 27.01 27.01 27.01 39.39 39.39 39.39 49.35 49.35 49.35 38.58
ClinicalCamel-70B 28.34 28.34 28.34 38.75 38.75 38.75 48.05 48.05 48.05 38.37
CancerLLM(Ours) 34.49 34.49 34.49 46.03 46.03 46.03 54.62 54.62 54.62 45.04
Table 3: Results of Cancer Diagnosis Generation on ICDdiagnosis
Exact Match BLEU-2 ROUGE-L F1
Approach Precision Recall F1 Precision Recall F1 Precision Recall F1 Average
PMC LLaMA 7B 88.61 88.61 88.61 90.36 90.36 90.36 93.26 93.26 93.26 90.74
Medalpaca 7B 89.28 89.28 89.28 91.27 91.27 91.27 93.89 93.89 93.89 91.48
LLAMA-2 7B 89.18 89.18 89.18 90.83 90.83 90.83 93.53 93.53 93.53 91.18
Mistral 1*7B 89.47 89.47 89.47 91.30 91.30 91.30 94.19 94.19 94.19 91.65
Mixtral 8*7B 90.23 90.23 90.23 92.05 92.05 92.05 94.50 94.50 94.50 92.26
Bio-Mistral 7B 88.90 88.90 88.90 91.05 91.05 91.05 93.79 93.79 93.79 91.24
LLama3 8B 89.94 89.94 89.94 91.75 91.75 91.75 94.34 94.34 94.34 92.01
MedLLaMA 13B 88.80 88.80 88.80 90.87 90.87 90.87 93.31 93.31 93.31 90.99
PMC LLaMA 13B 87.95 87.95 87.95 89.64 89.64 89.64 92.58 92.58 92.58 90.06
Medalpaca 13B 88.61 88.61 88.61 90.37 90.37 90.37 92.94 92.94 92.94 90.64
LLaMA2 13B 89.85 89.85 89.85 91.54 91.54 91.54 94.21 94.21 94.21 91.86
LLaMA2 70B 90.04 90.04 90.04 91.62 91.62 91.62 93.98 93.98 93.98 91.88
Llama3-OpenBioLLM-70B 88.33 88.33 88.33 90.02 90.02 90.02 93.15 93.15 93.15 90.50
ClinicalCamel-70B 92.02 92.02 92.02 93.62 93.62 93.62 95.52 95.52 95.52 93.72
CancerLLM(Ours) 89.37 89.37 89.37 91.98 91.98 91.98 93.98 93.98 93.98 91.78
Table 4: Results of Cancer Phenotype Extraction
3.1.2 Cancer Phenotype Extraction
Table 4 presents the experiment results of various approaches on the treatment plan generation task. We have the following observations: (1) Our CancerLLM significantly outperforms all strong baselines with the same parameter numbers or number of LLMs across all evaluation metrics. (2) Our CancerLLM performs comparably with LLMs that have larger parameters. For instance, LLama3 8B achieves 92.01%, and LLama2 13B achieves 91.86%. (3)CancerLLM could improve the performance of original Mistral 1*7B, and Bio-Mistral 7B in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (4) We observe that ClinicalCamel-70B achieves the best phenotype extraction performance; however, it has a larger number of parameters, which affects both training and inference time, as well as memory consumption. (5) Despite our CancerLLM not showing significant improvement across all LLMs with different parameters on the phenotype extraction task, as depicted in Figure 1, our model attains the best performance across the three tasks examined in this paper.
3.2 Results of Robustness Testbeds
3.2.1 Counterfactual Robustness
Exact Match BLEU-2 ROUGE-L
Model rate Precision Recall F1 Precision Recall F1 Precision Recall F1 Average F1
Mistral 1*7B 20% 88.01 88.01 88.01 89.20 89.20 89.20 94.10 94.10 94.10 72.53
40% 84.82 84.82 84.82 86.85 86.85 86.85 89.58 89.58 89.58
60% 77.13 77.13 77.13 79.49 79.49 79.49 82.66 82.66 82.66
80% 31.97 31.97 31.97 32.29 32.29 32.29 34.24 34.24 34.24
CancerLLM(Ours) 20% 88.05 88.05 88.05 89.90 89.90 89.90 92.26 92.26 92.26 78.03
40% 84.54 84.54 84.54 86.71 86.71 86.71 89.56 89.56 89.56
60% 77.51 77.51 77.51 79.18 79.18 79.18 82.26 82.26 82.26
80% 54.55 54.55 54.55 55.39 55.39 55.39 56.53 56.53 56.53
Table 5: Counterfactual robustness performance on phenotype extraction. The rate refers to different counterfactual rate.
To validate the robustness of our model, we propose a counterfactual robustness testbed designed to simulate incorrect annotations. Table 5 presents the complete results under various counterfactual rates. We observed that: (1) Our CancerLLM still achieves the best performance when compared to Mistrial 1*7B. (2) As the rate increases, the model’s performance deteriorates. (3) When the rate is set as 20%, 40%, and 60%, our model exhibits similar performance to Mistral 1*7B. However, when the rate is set at 80%, we observe that CancerLLM achieves higher F1 performance, indicating the robust counterfactual resilience of our model.
3.2.2 Misspellings Robustness
Exact Match BLEU-2 ROUGE-L
Model rate Precision Recall F1 Precision Recall F1 Precision Recall F1 Average F1
Diagnosis Generation Bio-Mistral 7B 2% 0.82 0.82 0.82 10.38 10.38 10.38 25.23 25.23 25.23 12.06
4% 0.64 0.64 0.64 11.78 11.78 11.78 27.01 27.01 27.01
6% 0.00 0.00 0.00 10.23 10.23 10.23 25.25 25.25 25.25
8% 0.09 0.09 0.09 9.56 9.56 9.56 23.80 23.80 23.80
CancerLLM(Ours) 2% 0.09 0.09 0.09 9.88 9.88 9.88 24.67 24.67 24.67 12.00
4% 0.00 0.00 0.00 10.29 10.29 10.29 25.37 25.37 25.37
6% 0.00 0.00 0.00 9.54 9.54 9.54 24.16 24.16 24.16
8% 0.27 0.27 0.27 10.61 10.61 10.61 25.81 25.81 25.81
ICDdiagnosis Bio-Mistral 7B 2% 0.00 0.00 0.00 7.07 7.07 7.07 18.00 18.00 18.00 8.62
4% 0.00 0.00 0.00 7.27 7.27 7.27 17.78 17.78 17.78
6% 0.27 0.27 0.27 8.07 8.07 8.07 18.63 18.63 18.63
8% 0.00 0.00 0.00 7.65 7.65 7.65 18.79 18.79 18.79
CancerLLM(Ours) 2% 0 0 0 8.93 8.93 8.93 20.59 20.59 20.59 9.57
4% 0.27 0.27 0.27 9.07 9.07 9.07 20.53 20.53 20.53
6% 0.27 0.27 0.27 8.34 8.34 8.34 19.34 19.34 19.34
8% 0 0 0 8.20 8.20 8.20 19.31 19.31 19.31
Table 6: Misspellings robustness performance on diagnosis generation and ICDdiagnosis. The rate refers to different misspelling rates in each instance.
To validate the robustness of our model, we propose a misspelling robustness testbed designed to simulate misspelling in the clinical notes. Table 6 presents the complete results and average F1 performance of different counterfactual rates across exact match, BLEU-2, and ROUGE-L metrics for both diagnosis generation and ICDdiagnosis. Our observations are as follows: CancerLLM consistently outperforms Bio-Mistral 7B. For the diagnosis generation task, CancerLLM achieves a similar average F1 score compared to Bio-Mistral 7B. Specifically, in ICD diagnosis, CancerLLM scores 9.57%, while Bio-Mistral 7B scores 8.62%.
3.3 Results of Generation Time and GPU Memory
Phenotype Extraction Diagnosis Generation1
LLMs F1 Time Used Memory F1 Time Used Memory
Bio-Mistral 7B 91.24 1:06:55 5,746 MB 68.89 1:07:45 5,802 MB
Mistral 1*7B 91.65 57:05 5,598 MB 55.07 1:07:49 5,680 MB
Mixtral 8*7B 92.26 2:01:27 25,086 MB 60.12 2:16:14 25,166 MB
PMC LLaMA 13B 90.06 1:08:59 8,208 MB 62,86 1:19:52 9,208 MB
LLaMA2 13B 91.86 1:08:43 8,204 MB 55.42 1:24:17 9,254 MB
ClinicalCamel-70B 93.72 2:50:16 37,716 MB 63.55 3:05:37 37,67 MB
CancerLLM(Ours) 91.78 1:14:12 5,550 MB 86.81 1:26:33 5,768 MB
Table 7: Comparation of Generation Time (hours:minutes:seconds) and Used GPU Memory (Used Memory: Megabyte (MB)) of different LLMs, F1 refers to the average F1
In Table 7, we compared the generation time on the whole testing set and used GPU memory on a single A100 across different LLMs. For the Phenotype Extraction task, we set the maximum input token length to 1500 and the maximum new token length to 50. In the Diagnosis Generation task, we set the maximum input token length to 1500 and the maximum new token length to 500. We observed that although Mixtral 8*7B, LLaMA2 13B, and ClinicalCamel-70B exhibit better phenotype extraction performance, they have higher generation time and GPU memory usage. For example, ClinicalCamel-70B requires 2:50:16 for inference and uses 37,716 MB of GPU memory, while CancerLLM requires only 1:14:12 for inference and uses just 5,550 MB of GPU memory.
4 Discussion
4.1 Cancer Diagnosis Generation
In this task, we primarily pre-train CancerLLM and explore its effectiveness in the diagnosis generation task by providing relevant information. As shown in Table 2, the Bio-Mistrial 7B obtains the best performance among all baseline LLMs. We guess that one reason for this is that Bio-Mistral 7B is trained on a large pre-processed corpus, such as the PubMed Central corpus Canese & Weis (2013). Another reason is that the foundational structure of Mixtral enhances training effectiveness. The Mixtral 8*7B outperforms the Mixtral 1*7B by more than 5% in average F1 value. However, Mixtral 8*7B is a Mixture of Experts (MoE) model with 8 experts, making training and inference time challenging. Fortunately, our CancerLLM achieves the best model performance through pre-training and fine-tuning with a single expert model. Not all LLMs with large parameters achieve the best performance. For example, the 70B models (LLaMA2 70B, Llama3-OpenBioLLM-70B, and ClinicalCamel-70B) do not outperform the Bio-Mistral 7B, despite being trained on medical corpora. This indicates that smaller parameter LLMs can achieve better performance. For ICDdiagnosis, we find that overall performance is lower than in diagnosis generation. This is because the expected diagnoses in ICD diagnosis are more complex. Our 7B model, CancerLLM, which incorporates cancer domain knowledge, also supports this point. This is particularly important in the medical domain as it reduces training time and memory usage, providing opportunities for medical institutions with limited computational resources.
4.2 Cancer Phenotype Extraction
In Table 4, We observed that the ClinicalCamel-70B model outperforms all baseline models in terms of average F1 value. However, its large parameter size significantly impacts both training and inference times, making it less efficient and more resource-intensive. On the other hand, our 7B model, CancerLLM, shows promising results, performing comparably to both the Mixture of Experts (MoE) model Mistrial 8*7B and the LLamA2 13B model. This is particularly noteworthy given CancerLLM’s smaller parameter size, which makes it more efficient. We suspect that the reason behind CancerLLM’s performance lies in the specific nature of the clinical notes and pathology reports we used for training. These documents may not contain sufficient annotation information, which is crucial for extracting phenotypic data. This highlights the importance of data quality and annotation in training effective LLM models for clinical information extraction tasks.
4.3 Robustness
The proposed two testbeds were primarily used to evaluate the counterfactual robustness and misspelling robustness of the LLMs. In the phenotype extraction task, when the counterfactual rate is set to 20%, we observed that the performance of CancerLLM and Mixtrual 1*7B does not significantly decrease. However, when the rate exceeds 60%, there is a notable decline in F1 performance. This indicates that Mixtrual 1*7B and CancerLLM can handle scenarios with fewer counterfactual instances. However, we found that when varying the misspelling rate, both the Bio-Mistral 7B and CancerLLM models experienced a significant decline in performance. This is particularly evident in the exact match metric, where performance drops close to zero. These results highlight the critical importance of correct spelling for LLMs to generate accurate diagnoses and treatment plans, as even minor misspellings can severely impact the models’ ability to correctly interpret and process medical information. This underscores the necessity for meticulous data preprocessing and validation in clinical settings to ensure that the input data is free of errors, thereby enabling the models to function at their highest potential and provide reliable outcomes.
4.4 Error Analysis
4.4.1 Cancer Diagnosis Generation
Error type Expected output Error output
Incomplete Generation {breast and lung cancer} {metastatic breast cancer}
{nsclc t4n2mo-1a} {nsclc t4n2mo}
{metastatic lung cancer with brain mets} {metastatic lung cancer}
Irrelevant Generation { dcis (Ductal carcinoma in situ) } {invasive ductal carcinoma}
{nsclc} { lung cancer}
Misspelling { metastatic lung cnacer } {metastatic lung cancer}
Redundant Generation { nsclc} {nsclc,t2n0, s/p lobectomy and chemotherapy}
{lung cancer} { lung cancer with brain mets}
Abbreviation { lung ca } {lung cancer}
Table 8: Error generation cases of diagnosis generation. Note: Due to privacy reasons, we did not provide the relevant clinical notes for each error type.
As shown in Table 8, we have comprehensively summarized two primary instances of error generation observed in CancerLLM. Our observations reveal the following:
  • (1) Misspellings and abbreviations within the clinical notes exert a significant influence on both the training and evaluation phases of the model. These linguistic inaccuracies can mislead the model’s learning process and hinder its ability to accurately generate diagnoses.
  • (2) Another notable observation is the model’s tendency to produce incomplete output in certain scenarios. For instance, when presented with a ground diagnosis such as metastatic lung cancer with brain metastases, the model may overlook crucial details such as the presence of brain metastases, This omission diminishes the comprehensiveness and accuracy of the generated diagnoses, potentially leading to misinterpretations and suboptimal clinical decisions.
These findings underscore the importance of meticulous data preprocessing and model refinement to mitigate the impact of linguistic nuances and enhance the CancerLLM’s diagnostic capabilities.
4.4.2 Cancer Phenotype Extraction
Error type Input sentence Expected output Error output
Redundant Generation the cores are infiltrated by ductal carcinoma growing in nests and sheets with areas of residual lumen formation what is the histological type in the given context? {ductal carcinoma} {ductal carcinoma, residual lumen formation}
Irrelevant Generation one requisition slip and labeled right breast 1:00, part b on the container, please describe the tumor location in the given context {1:00} {it is not a relevant question}
Repeat Generation In the text ypt2 ypn1bii mo, stage iib, what is the stage of cancer in the given context? {ypn1bii, iib,mo,ypt2} {iib,iib,iib,iib,iib, iib,iib,iib}
Incomplete Generation In the text er/pr+, her2 negative, please identify the receptors mentioned in the provided context {pr, her2, er} {pr,er}
In the text sections from the stellate area in the medical breast at the 9-10:00 position show invasive, please describe the tumor location. {medical breast at the 9-10:00} {9-10:00 position}
Inaccurate Annotation In the text ductal carcinoma, solid, intermediate to high nuclear grade 3, what is the grade of cancer? {high, intermediate} {3}
Table 9: Error generation cases of cancer phenotype extraction
As shown in Table 9, we have identified and summarized six primary error-generation cases, shedding light on crucial aspects influencing the model’s performance. Our observations are as follows:
  • Redundant Generation: One prevalent issue is the model’s tendency to generate repetitive and incomplete outputs when confronted with insufficient contextual information. This limitation hampers the model’s ability to provide comprehensive and accurate responses.
  • Abbreviation Challenges: Abbreviations pose a significant challenge for CancerLLM, as evidenced by instances such as those outlined in Table 9 (Incomplete generation), where abbreviations like pr and er lead to incomplete outputs. Resolving this challenge is vital for enhancing the model’s interpretive accuracy.
  • Contextual Misinterpretation: CancerLLM struggles to comprehend questions accurately within the given context, as illustrated in Table 9 (Redundant generation), where it misidentifies residual lumen formation as a histological type. This highlights the need for improved contextual understanding to prevent such errors.
  • Annotation Quality: The quality of annotation data for phenotype extraction significantly impacts CancerLLM’s learning process. Inaccurate annotations, as discussed, can impede the model’s ability to learn effectively, underscoring the importance of providing high-quality annotation data to facilitate robust learning outcomes.
These findings emphasize the multifaceted nature of the challenges faced by CancerLLM and underscore the importance of addressing these issues to enhance its performance and reliability in clinical applications.
5 Conclusion
In this study, we introduced CancerLLM, a medical LLM designed specifically for the cancer domain. We provided three datasets for cancer phenotype extraction, cancer diagnosis generation. Additionally, we proposed two testbeds to evaluate the robustness of CancerLLM. Comparing CancerLLM with 14 widely used LLMs, our results demonstrate that CancerLLM achieves state-of-the-art performance across three tasks. Our work provides invaluable insights and tools for further research on leveraging AI to enhance cancer domain.
6 Acknowledgments
This research was supported by the National Institutes of Health’s National Center for Complementary and Integrative Health grant number R01AT009457, National Institute on Aging grant number R01AG078154 and National Cancer Institute grant number R01CA287413. The content is solely the responsibility of the authors and does not represent the official views of the National Institutes of Health. We thank support from UMN’s Center for Leanring Health System Sciences. We thank support from UMN Data Science Initiative seed grant. This research was supported by the National Institutes of Health’s National Center for Advancing Translational Sciences, grant UM1TR004405. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health’s National Center for Advancing Translational Sciences.
7 Competing Interests
The authors declare no competing financial or non-financial interests.